# Task Arithmetic Through The Lens Of One-Shot Federated Learning

### v1

@@ 3.2 Training Heterogeneity
This subsection examines the effect of different task objective function $L_{t}$ being optimized with varying local learning rates and numbers of iterations during the local training, which we refer to as training heterogeneity.
In the previous section, we assumed that all task objective functions are optimized using a homogeneous training process with a fixed number of iterations $K$ and constant learning rate $\eta$. However, in practice, each task objective function is often optimized with different hyperparameter settings, which introduces training heterogeneity and can lead to objective inconsistency [69]. Now, we extend the setting in Section 3.1 to consider each task objective function $L_{t}$ optimized with distinct hyperparameters $\eta_{t}^{(k)}$ and $K_{t}$. We adopt notation and apply theoretical insights from [69] to illustrate the impact of local training heterogeneity.
First, we define the following matrix of stochastic gradients for each task $t$
$G_{t}=[g_{t}(\theta_{t}^{(0)})\quad g_{t}(\theta_{t}^{(1)})\quad\dots\quad g_{t}(\theta_{t}^{(K_{t}-1)})]\in\mathbb{R}^{d\times K_{t}}$
where $g_{t}$ is the stochastic gradient of $L_{t}$. Next we define the following vector of normalized learning rates for each task $t$ as
$a_{t}=\begin{bmatrix}\frac{\eta_{t}^{(0)}}{\eta}\\ \frac{\eta_{t}^{(1)}}{\eta}\\ \vdots\\ \frac{\eta_{t}^{(K_{t}-1)}}{\eta}\end{bmatrix}\in\mathbb{R}^{K_{t}}$
@@
where the second equality follows from $\lambda=\frac{\beta}{T}$ as mentioned in Section 2. Next, we denote $\tau_{\text{eff}}=\frac{\beta}{T}\sum_{t=1}^{T}\|a_{t}\|_{1}$ as the effective number of steps which measures the average amount of updates accumulated using the constant learning rate $\eta$, and denote $w_{t}=\frac{\|a_{t}\|_{1}}{\sum_{s=1}^{T}\|a_{s}\|_{1}}$ as the aggregation weight for task $t$. Then we can further rewrite equation (8) as
@@
Notice the weight coefficients vector $[w_{1};w_{2};\dots;w_{T}]$ differs from the original uniform coefficients $[\frac{1}{T};\frac{1}{T};\dots;\frac{1}{T}]$ in the objective function $L$ (equation (1)). This is a discrepancy caused by training heterogeneity. In fact, the discrepancy between $[w_{1}\;w_{2}\;\dots\;w_{T}]$ and $[\frac{1}{T}\;\frac{1}{T}\;\dots\;\frac{1}{T}]$ leads FedAvg with multiple communication rounds to converge to the stationary point of a different objective function
$\tilde{L}(\theta):=\sum_{t=1}^{T}w_{t}L_{t}(\theta)$
which is inconsistent with the original objective function $L$. While Task Arithmetic involves only a single round of FedAvg, the inconsistency still remains due to training heterogeneity. Formally, we present the following assumptions and adapt Theorems 1 and 2 from [69] to contextualize this inconsistency in our setting.
@@
(Smoothness) Assume all the task objective functions $L_{t}$ are $H$-smooth. That is, $\forall t\in[T]$ and $\forall\theta,\varphi\in\mathbb{R}^{d}$,
$\|\nabla L_{t}(\theta)-\nabla L_{t}(\varphi)\|\leq H\|\theta-\varphi\|.$
@@
Notice that Assumption 3.8 imposes a more restrictive condition on data heterogeneity compared to Assumption 3.4. Currently, no unified notion of data heterogeneity exists for Federated Learning. Since this section focuses on training heterogeneity, we adopt this more restrictive notion of data heterogeneity, as done in [69], to facilitate theoretical development.
@@
(Theorem 1 and 2 from [69]) Consider $\theta_{TA}$ from update rule (8). Denote $\tilde{L}(\theta)=\sum_{t=1}^{T}w_{t}L_{t}(\theta)$ and $\bar{K}=\frac{1}{T}\sum_{t=1}^{T}K_{t}$. Let $\eta=\sqrt{T/\bar{K}}$. Under Assumption 3.2,3.7 and 3.8, we have the following bound on the gradient norm $\|\nabla\tilde{L}(\theta_{TA})\|^{2}$:
@@
The theorem above illustrates the impact of a heterogeneous local training process on Task Arithmetic. When different training processes are used for each objective function, the chi-square divergence between the weight coefficient vectors becomes non-zero, resulting in a persistent error term, $\chi^{2}_{p||w}\zeta^{2}$. This error term only vanishes if $\zeta^{2}=0$, indicating minimal data heterogeneity. This relationship highlights the interaction between data heterogeneity and training heterogeneity: significant data heterogeneity exacerbates the negative effects of training heterogeneity, intensifying the overall performance degradation.
When all task objective functions are optimized with the same number of iterations $K$ and a consistent learning rate schedule $\{\eta^{(0)},\dots,\eta^{(K-1)}\}$, we have $w_{t}=\frac{1}{T}$. This yields $\chi_{p||w}^{2}=0$, aligning the actual objective function $\tilde{L}$ being optimized with the original objective function $L$. In this scenario, objective inconsistency is effectively eliminated, as the tasks are uniformly weighted and trained under identical conditions.
@@ 4 Adapting Federated Learning Algorithms for Task Arithmetic
In the previous section, we used insights from FedAvg to analyze how data and training heterogeneity negatively impact Task Arithmetic. In order to address these challenges, numerous algorithms have been developed to improve FedAvg for more efficient Federated Learning (see Section 7 for details).
Thus, we adapt some Federated Learning algorithms for Task Arithmetic to solve heterogeneity challenges for better model merging performance. To guide this adaptation, we establish specific criteria for selecting suitable algorithms. Additional motivations and challenges for this adaptation are discussed in Appendix B.
  • Adaptability to One-Shot Setting: Algorithms must be effective in a single communication round since multiple rounds are infeasible in Task Arithmetic.
  • No Additional Training Required: Algorithms that significantly increase computational costs to address heterogeneity are unsuitable, as Task Arithmetic’s key advantage is its minimal computational overhead.
  • No Access to Additional Datasets Required: Algorithms relying on external datasets, such as those used in knowledge distillation, are impractical due to data constraints.
@@
With these criteria established, we now explore four Federated Learning algorithms FedNova [69], FedGMA [66], Median [85] and CCLIP [33] that align with these guidelines, explaining their motivations and how they modify Task Arithmetic. For more detailed explanation and further understanding of these algorithms, please refer to the original papers.
4.1 FedNova [69]
@@
FedNova addresses objective inconsistency caused by training heterogeneity by replacing the heterogeneous weight vector $[w_{1}\;w_{2}\;\dots\;w_{T}]$ with the uniform weight vector $[\frac{1}{T}\;\frac{1}{T}\;\dots\;\frac{1}{T}]$, ensuring consistent weighting across tasks. This approach adapts easily to a one-shot setting. Using the notation from Section 3.2, FedNova modifies the Task Arithmetic update as:
@@
4.2 FedGMA [66]
FedGMA addresses data heterogeneity by mitigating sign conflicts among local updates, which in FedAvg can cause information loss and slower convergence. It achieves this by using a gradient mask to reduce the impact of conflicting directions and preserve meaningful information.
Specifically, FedGMA computes an agreement score $A$ to measure alignment across task vectors $\{\tau_{t}\}_{t=1}^{T}\subset\mathbb{R}^{d}$. Based on a threshold $\rho$, FedGMA constructs a mask $\tilde{M}$ that emphasizes coordinates with strong agreement while reducing the influence of others. Formally:
$A=\bigg{|}\frac{1}{T}\sum_{t=1}^{T}\operatorname{sign}(\tau_{t})\bigg{|}\quad\text{and}\quad\tilde{M}_{j}=\begin{cases}1,&\text{if }A_{j}\geq\rho\\ A_{j},&\text{otherwise}\end{cases}$
where $j$ denotes the $j$-th coordinate and $\operatorname{sign}(\cdot)$ is applied in a coordinate-wise manner. This yields
@@ 4.2 FedGMA [66]
4.3 Median [85]
Coordinate-Wise Median [85], originally designed to handle adversarial updates in Federated Learning, is adapted here to address data and training heterogeneity in Task Arithmetic. Due to diverse data distributions or differing hyperparameter settings, some task vectors may have extreme values. By selecting the median value for each coordinate, this method reduces the influence of outliers while maintaining overall performance across tasks. It modifies Task Arithmetic as
@@ 4.3 Median [85]
4.4 CCLIP [33]
CCLIP, short for centered clipping, is another widely applied robust aggregation method towards adversarial devices in Federated Learning. With the same motivation to using the Coordinate-Wise Median, we use CCLIP to reduce the impact of extreme task vectors. CCLIP is implemented with a predefined threshold $\rho$ and modifies Task Arithmetic as follows:
@@ 5 Experiments on Merging CLIP
In this section, we present and discuss our experimental results on CLIP-ViT-B-32 [58] for image classifications. We follow the same experimental paradigm as [28]. Specifically, we use CLIP-ViT-B-32 [58] as the pre-trained model and eight datasets: Cars [37], DTD [9], EuroSAT [26], GTSRB [62], MNIST [38], RESISC45 [8], SUN397 [78], and SVHN [53], to construct eight task vectors.
In total, there are 247 ways to select $T$ different task vectors from these eight task vectors where $T\in[2,8]$: $247=\sum_{T=2}^{8}{8\choose T}$. For each algorithm, we therefore conduct 247 experiments. In each experiment, we merge $T$ selected task vectors and evaluate on the $T$ datasets corresponding to the task vectors used. Our evaluation metric is normalized accuracy [28], defined as the test accuracy normalized by the fine-tuned model’s accuracy. That is,
@@ 5.1 Experimental Results
In the first part of the experiments (Section 5.1.1), to simulate practical conditions of training heterogeneity, we fine-tune CLIP-ViT-B-32 on each dataset using three learning rates $\{1\mathrm{e}{-4},1\mathrm{e}{-5},1\mathrm{e}{-6}\}$ and different numbers of iterations. Then we select the best fine-tuned checkpoints via cross-validation on validation datasets. Refer to Appendix C.1 for further details on fine-tuning and cross-validation.
In the second part of the experiments (Section 5.1.2), in order to better understand the impact of training heterogeneity, we use the task vectors provided by [28], which were fine-tuned with uniform training conditions—same number of iterations and learning rates—thereby eliminating training heterogeneity.
@@ Methods
Table 1: Combining all eight task vectors using five different methods. Each method is evaluated on eight datasets, with normalized accuracy reported for each. The highest and second-highest normalized accuracy values for each dataset are highlighted in bold and underlined, respectively.
@@
Percentage of
Improved Experiments
Percentage of
Unchanged Experiments
Percentage of
Degraded Experiments
@@
Figure 1: Histograms showing the change in average normalized accuracy for four different methods compared to Task Arithmetic. For each plot, the x-axis represents the change in average normalized accuracy, calculated as the difference between the average normalized accuracy of the algorithm used and that of Task Arithmetic. The y-axis indicates the number of experiments within the range of change values. A positive value on the x-axis indicates that the algorithm improves upon Task Arithmetic, while a negative value indicates that the algorithm degrades Task Arithmetic.
Table 2 and Figure 1 summarize the performance comparison between Task Arithmetic and other Federated Learning algorithms across 247 experiments. In Table 2, we report the percentage of 247 experiments in which the average normalized accuracy improves, remains unchanged, or degrades when using four Federated Learning methods compared to the baseline, Task Arithmetic. The average normalized accuracy is calculated by averaging over the number of task vectors being used. In order to better visualize the performance differences for each method, in Figure 1 we use histograms to show the frequencies of experiments within each range of change in average normalized accuracy. Median, FedNova, and CCLIP consistently show the ability to improve upon Task Arithmetic in most cases, while FedGMA typically demonstrates either no change or slight improvements to Task Arithmetic’s performance. Once again, we observe that Median and CCLIP exhibit the most significant improvements over Task Arithmetic.

### v2

@@ 4.2 Training Heterogeneity
This subsection examines the effect of training heterogeneity on Task Arithmetic, which we define as follows:
Definition 4.8.
(Training Heterogeneity) In Task Arithmetic, training heterogeneity refers to the phenomenon where different task objective functions $L_{t}$ are optimized with varying local learning rates and numbers of iterations during local training.
In the previous section, we assumed that all task objective functions are optimized with a fixed number of iterations $K$ and constant learning rate $\eta$. However, in practice, each task objective function is often optimized with different hyperparameter settings, which introduces training heterogeneity and can lead to objective inconsistency (Wang et al., 2020). Now, we extend the setting in Section 4.1 to consider each task objective function $L_{t}$ being optimized with distinct hyperparameters $\eta_{t}^{(k)}$ and $K_{t}$. We adopt notation and apply theoretical insights from Wang et al. (2020) to illustrate the impact of training heterogeneity.
First, let $g_{t}$ be the stochastic gradient of $L_{t}$, and we define the following matrix of stochastic gradients for each task $t$
$G_{t}=[g_{t}(\theta_{t}^{(0)})\quad g_{t}(\theta_{t}^{(1)})\quad\dots\quad g_{t}(\theta_{t}^{(K_{t}-1)})]\in\mathbb{R}^{d\times K_{t}}.$
Next we define the vector of normalized learning rates for each task $t$ as
$a_{t}=[\frac{\eta_{t}^{(0)}}{\eta},\frac{\eta_{t}^{(1)}}{\eta},\dots,\frac{\eta_{t}^{(K_{t}-1)}}{\eta}]^{T}\in\mathbb{R}^{K_{t}}$
@@
where the second equality follows from $\lambda=\frac{\beta}{T}$ as mentioned in Section 3. Next, we denote $\tau_{\text{eff}}=\frac{\beta}{T}\sum_{t=1}^{T}\|a_{t}\|_{1}$ as the effective number of steps which measures the average number of updates accumulated using the constant learning rate $\eta$, and denote $w_{t}=\frac{\|a_{t}\|_{1}}{\sum_{s=1}^{T}\|a_{s}\|_{1}}$ as the aggregation weight for task $t$. Then we can further rewrite equation (7) as
@@
Notice the weight coefficients vector $[w_{1};w_{2};\dots;w_{T}]$ differs from the original uniform coefficients $[\frac{1}{T};\frac{1}{T};\dots;\frac{1}{T}]$ in the objective function $L$ from equation (1). This discrepancy is caused by training heterogeneity, and leads FedAvg with multiple communication rounds to converge to the stationary point of a different objective function $\tilde{L}(\theta):=\sum_{t=1}^{T}w_{t}L_{t}(\theta)$ which is inconsistent with the original objective function $L$. While Task Arithmetic involves only one round of FedAvg, the inconsistency still remains due to training heterogeneity. Formally, we present the following assumptions and adapt Theorems 1 and 2 from Wang et al. (2020) to illustrate this inconsistency in our setting.
@@
(Smoothness) Assume all the task objective functions $L_{t}$ are $H$-smooth. That is, $\forall t\in[T]$ and $\forall\theta,\varphi\in\mathbb{R}^{d}$, $\|\nabla L_{t}(\theta)-\nabla L_{t}(\varphi)\|\leq H\|\theta-\varphi\|.$
@@
Notice that Assumption 4.10 imposes a more restrictive condition on data heterogeneity compared to Assumption 4.5. Currently, no unified notion of data heterogeneity exists for Federated Learning. Since this section focuses on training heterogeneity, we adopt this more restrictive notion of data heterogeneity, as done in Wang et al. (2020), to facilitate theoretical development.
@@
(Theorem 1 and 2 from (Wang et al., 2020)) Consider $\theta_{TA}$ from update rule (7). Denote $\tilde{L}(\theta)=\sum_{t=1}^{T}w_{t}L_{t}(\theta)$ and $\bar{K}=\frac{1}{T}\sum_{t=1}^{T}K_{t}$. Let $\eta=\sqrt{T/\bar{K}}$. Under Assumption 4.3,4.9 and 4.10, we have the following bound on the gradient norm $\|\nabla\tilde{L}(\theta_{TA})\|^{2}$:
@@
The theorem above illustrates how heterogeneous local training affects the convergence of Task Arithmetic to a stationary point of $L$. The convergence rate in (10) is influenced by two key factors: the term $\chi^{2}_{p||w}\zeta^{2}$ and the convergence rate to a stationary point of $\tilde{L}$ in (9). For $\chi^{2}_{p||w}\zeta^{2}$, when different training processes are used for each objective function, $\chi^{2}_{p||w}$ is non-zero, resulting in $\chi^{2}_{p||w}\zeta^{2}$ as a persistent error term. With training heterogeneity, $\chi^{2}_{p||w}\zeta^{2}$ only vanishes if $\zeta^{2}=0$, indicating minimal data heterogeneity. This highlights the interaction between data heterogeneity and training heterogeneity: significant data heterogeneity exacerbates the negative effects of training heterogeneity, intensifying the overall performance degradation.
When all task objective functions are optimized with the same number of iterations $K$ and a consistent learning rate schedule $\{\eta^{(0)},\dots,\eta^{(K-1)}\}$, we have $w_{t}=\frac{1}{T}$. This yields $\chi_{p||w}^{2}=0$ and $L=\tilde{L}$, aligning the actual objective function $\tilde{L}$ being optimized with the original objective function $L$. In this scenario, objective inconsistency is effectively eliminated. The convergence rate in (10) is reduced to (9).
@@ 5 Adapting Federated Learning Algorithms for Task Arithmetic
In the previous section, we used insights from FedAvg to analyze how data and training heterogeneity negatively impact Task Arithmetic. In order to address these challenges, numerous algorithms have been developed to improve FedAvg for more efficient Federated Learning (see Section 2 for related work). Thus, we adapt some Federated Learning algorithms to solve heterogeneity challenges in Task Arithmetic for better model merging performance. Selecting the right Federated Learning algorithms to implement requires a clear understanding of key challenges that complicate the adaptation. In this section, we begin with analyzing several key challenges.
5.1 Challenges in Adapting Federated Learning Algorithms for Task Arithmetic
First, the number of communication rounds is limited. Since Task Arithmetic is only one-shot Federated Learning, algorithms relying on multiple communication rounds are unsuitable. For instance, some Federated Learning algorithms add regularization terms to local objective functions (Li et al., 2020b; Durmus et al., 2021) to encourage local updates to remain close to the global model parameters transmitted from the previous communication round. However, in our one-shot setting, only the pre-trained model parameters $\theta_{0}$ are communicated. Applying such regularization would force each task’s fine-tuned parameters to remain close to $\theta_{0}$, potentially hindering both convergence and task-specific performance. Similarly, algorithms that address heterogeneity through iterative techniques such as variance reduction (Karimireddy et al., 2020; Reddi et al., 2020) or adaptive step size tuning (Jhunjhunwala et al., 2023; Li et al., 2023b) require the accumulation and updating of certain metrics over multiple communication rounds. Since Task Arithmetic operates within a one-shot setting, implementing such iterative updates is impossible.
Second, no additional training is allowed. To counteract the effects of heterogeneity or limited communication and data budget, some Federated Learning algorithms require additional training, imposing additional computational cost. For instance, Ye et al. (2023b) ask each device to learn metrics that compare local and global data through training, and these metrics are then used as additional scores for aggregation. Zhang et al. (2022) train an additional generator to generate data which is subsequently used to train a global model in one-shot Federated Learning setting.
Third, no additional datasets are available. Many Federated Learning algorithms rely on supplementary datasets, which is not feasible for modifying Task Arithmetic. In one-shot Federated Learning, a common approach to address data heterogeneity is knowledge distillation (Zhou et al., 2020; Guha et al., 2019; Li et al., 2020a). These methods often require access to extra datasets from which either local devices or the central server distills knowledge to improve model performance.
Given the constraints and unique needs for adapting Federated Learning algorithms, we identify the following three criteria for selecting Federated Learning algorithms: adaptability to one-shot setting, no additional training required and no access to additional datasets required.
5.2 Algorithms
@@
With criteria established above, we now explore four Federated Learning algorithms FedNova (Wang et al., 2020), FedGMA (Tenison et al., 2022), Median (Yin et al., 2018) and CCLIP (Karimireddy et al., 2021) that align with these guidelines, explaining their motivations and how they modify Task Arithmetic. For more detailed explanation and further understanding of these algorithms, please refer to the original papers.
@@
FedNova (Wang et al., 2020) FedNova addresses objective inconsistency caused by training heterogeneity by replacing the heterogeneous weight vector $[w_{1}\;w_{2}\;\dots\;w_{T}]$ with the uniform weight vector $[\frac{1}{T}\;\frac{1}{T}\;\dots\;\frac{1}{T}]$, ensuring consistent weighting across tasks. This approach adapts easily to the one-shot setting. Using the notation from Section 4.2, FedNova modifies the Task Arithmetic update as:
@@
FedGMA (Tenison et al., 2022) FedGMA addresses data heterogeneity by mitigating sign conflicts among local updates. Such sign conflicts can cause information loss and slower convergence. FedGMA uses a gradient mask to reduce the impact of conflicting directions and preserves meaningful information. Specifically, it computes an agreement score $A$ to measure alignment across task vectors $\{\tau_{t}\}_{t=1}^{T}$. Based on a threshold $\rho$, FedGMA constructs a mask $\tilde{M}$ that emphasizes coordinates with strong agreement while reducing the influence of others. Formally, define $A=\bigg{|}\frac{1}{T}\sum_{t=1}^{T}\operatorname{sign}(\tau_{t})\bigg{|}$ and $\tilde{M}_{j}=\begin{cases}1,&\text{if }A_{j}\geq\rho\\ A_{j},&\text{otherwise}\end{cases}$ where $j$ denotes the $j$-th coordinate and $\operatorname{sign}(\cdot)$ is applied in a coordinate-wise manner. This yields
@@
Median (Yin et al., 2018) Coordinate-Wise Median (Yin et al., 2018), originally designed to handle adversarial updates in Federated Learning, is adapted here to address data and training heterogeneity in Task Arithmetic. Due to diverse data distributions or differing hyperparameter settings, some task vectors may have extreme values. By selecting the median value for each coordinate, this method reduces the influence of outliers while maintaining overall performance across tasks. It modifies Task Arithmetic as
@@
CCLIP (Karimireddy et al., 2021) CCLIP, short for centered clipping, is another widely applied robust aggregation method towards adversarial devices in Federated Learning. With the same motivation to using the Coordinate-Wise Median, we use CCLIP to reduce the impact of extreme task vectors. CCLIP is implemented with a predefined threshold $\rho$ and modifies Task Arithmetic as follows:
@@ 6 Experiments on Merging CLIP
In this section, we present and discuss our experimental results on CLIP-ViT-B-32 (Radford et al., 2021) for image classifications. We follow the same experimental paradigm as (Ilharco et al., 2023). Specifically, we use CLIP-ViT-B-32 (Radford et al., 2021) as the pre-trained model and eight datasets: Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), RESISC45 (Cheng et al., 2017), SUN397 (Xiao et al., 2016), and SVHN (Netzer et al., 2011), to construct eight task vectors.
In total, there are 247 ways to select $T$ different task vectors from these eight task vectors where $T\in[2,8]$: $247=\sum_{T=2}^{8}\binom{8}{T}$. For each algorithm, we therefore conduct 247 experiments. In each experiment, we merge $T$ selected task vectors and evaluate on the $T$ datasets corresponding to the task vectors used. Our evaluation metric is normalized accuracy (Ilharco et al., 2023), defined as the test accuracy normalized by the fine-tuned model’s accuracy. That is,
@@ 6.1 Experimental Results
In the first part of the experiments (Section 6.1.1), to simulate practical conditions of training heterogeneity, we fine-tune CLIP-ViT-B-32 on each dataset using three learning rates $\{1\mathrm{e}{-4},1\mathrm{e}{-5},1\mathrm{e}{-6}\}$ and different numbers of iterations. Then we select the best fine-tuned checkpoints via cross-validation on validation datasets. Refer to Appendix B.1 for further details on fine-tuning and cross-validation.
In the second part of the experiments (Section 6.1.2), to better understand the impact of training heterogeneity, we use the task vectors provided by Ilharco et al. (2023), which were fine-tuned with the same number of iterations and the same learning rates, thereby eliminating training heterogeneity.
@@ Methods
Table 1: Merging all eight task vectors using five different methods. Each method is evaluated on eight datasets, with normalized accuracy reported for each dataset. The highest and second-highest normalized accuracy values for each dataset are highlighted in bold and underlined, respectively.
@@
Percentage of Improved Experiments
Compared to Task Arithmetic
Percentage of Unchanged Experiments
Compared to Task Arithmetic
Percentage of Degraded Experiments
Compared to Task Arithmetic
@@
Figure 1: Histograms showing the change in average normalized accuracy for four different methods compared to Task Arithmetic. The x-axis shows the difference in average normalized accuracy relative to Task Arithmetic, with positive values indicating improvement and negative values indicating degradation. The y-axis represents the number of experiments within each range of change values.
Table 2 and Figure 1 summarize the performance comparison between Task Arithmetic and other Federated Learning algorithms across 247 experiments. In Table 2, we report the percentage of 247 experiments in which the average normalized accuracy improves, remains unchanged, or degrades when using four Federated Learning methods compared to the baseline, Task Arithmetic. The average normalized accuracy is calculated by averaging over the number of task vectors being used. In order to better visualize the performance differences for each method, in Figure 1, we use histograms to show the frequencies of experiments within each range of change in average normalized accuracy. Median, FedNova, and CCLIP consistently show the ability to improve upon Task Arithmetic in most cases, while FedGMA typically demonstrates either no change or slight improvements.